OpenStudy (anonymous):

2(j+k)=m Solve for k

OpenStudy (anonymous):

2(j+k)=m 2j+2k=m 2k=m-2j k=m/2 -2j

OpenStudy (anonymous):

i get the shown work but u lost me at the solution?